OpenStudy (anonymous):

Helllp :3 Directions: Suppose y varies directly as x. Write a direct variation equation that relates x any y. The solve. 1.If y=-11 when x=6, find x when y=44

OpenStudy (anonymous):

\[y \space \alpha \space x\] \[y \space = k \space x \space , \space where \space k \space is \space a \space constant\] \[k \space = \space y \space \div x \space \] \[y = -11 \space when \space x = -6 \space , \space \space k = -6 \div -11 = \] \[y = \frac{ 6 }{ 11 } x\] when y=44, \[x = \frac{ 11 }{ 6 } \times 44 = 24 \]